# Enhancing Zero-Shot Facial Expression Recognition by LLM Knowledge Transfer

### v1

@@ IV-D Ablation Analysis
Evaluation of the Projection Head Effectiveness: Our method is built on the CLIP model [20], to evaluate the effectiveness of the learned projection head, we compare our method to the baseline CLIP model from two views: various ViT models and various zero-shot prompts. Our method uses the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”. As shown in Tab. II, under different ViT models, our method outperforms the plain CLIP model on almost all the test sets (except one UAR on FERPlus and one WAR on DFER). With ViT-B-32, our method improves the average WAR by 1.67% and average UAR by 6.14% over eight test sets. With ViT-B-16, our method improves the average WAR by 0.38% and average UAR by 4.56% over eight test sets. With ViT-L-14, our method improves the average WAR by 5.85% and average UAR by 13.67% over eight test sets. Furthermore, we compare our method with CLIP using different prompts, in which both single prompt and prompt ensembles are considered. Specifically, for a single prompt, we utilize “{class name}.”, “an expression of {class name}.”, and “a photo of a face with an expression of {class name}.”, respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations and the ten text prompt templates are outlined below:
TABLE V: Zero-shot FER results of our method with different instructions. Both task-related and task-unrelated instructions are investigated. All models use the prompt “an expression of [class].” for zero-shot prediction.
@@
ViT-B-32
Task Unrelated 40.49/47.62 31.93/31.92 27.93/27.93 34.95/43.32 25.55/22.05 24.06/20.10 18.38/21.66 31.54/31.50
42.08/42.57 33.86/33.87 29.98/29.98 42.99/39.24 24.42/20.95 21.78/18.11 18.01/21.09 32.22/31.76
38.12/36.05 31.81/31.81 27.53/27.53 37.89/35.07 25.84/23.15 20.23/16.78 16.85/19.32 31.08/29.92
Mean 40.23/42.08 32.53/32.53 28.48/28.48 38.61/39.21 25.27/22.05 22.02/18.33 17.75/20.69 31.61/31.06
Variance 1.99/16.82 0.66/0.67 0.86/0.86 8.27/8.51 0.28/0.61 1.86/1.40 0.32/0.74 0.16/0.5
Task Related 41.38/49.35 33.07/33.07 29.16/29.16 39.65/45.04 28.09/25.98 23.01/20.26 19.62/22.71 31.34/31.50
41.60/49.45 33.24/33.24 29.03/29.03 39.64/45.04 28.01/25.88 22.89/20.17 19.67/22.78 31.88/32.02
41.77/49.67 32.95/32.95 28.83/28.83 40.09/44.34 28.34/26.47 22.96/20.47 19.83/23.12 31.25/31.50
Mean 41.58/49.49 33.09/33.09 29.01/29.01 39.79/44.81 28.15/26.11 22.95/20.3 19.71/22.87 31.49/31.67
Variance 0.02/0.01 0.01/0.01 0.01/0.01 0.03/0.08 0.01/0.05 0.00/0.01 0.01/0.02 0.06/0.05
ViT-B-16
Task Unrelated 51.79/48.99 37.55/37.55 32.85/32.86 41.94/48.49 28.96/24.73 24.43/21.88 19.27/21.42 34.08/33.33
53.45/52.35 41.61/41.61 35.48/35.48 42.87/44.47 28.53/27.12 24.86/24.37 20.26/22.47 33.89/33.60
46.04/41.59 36.20/36.21 31.63/31.63 36.12/44.34 24.11/21.07 20.26/18.55 16.54/18.59 27.12/26.51
Mean 50.43/47.64 38.45/38.46 33.32/33.32 40.31/45.77 27.20/24.31 23.18/21.60 18.69/20.83 31.70/31.15
Variance 7.56/15.15 3.96/3.95 1.94/1.93 6.69/2.78 3.60/4.64 3.23/4.26 1.86/2.01 7.86/8.07
Task Related 56.25/58.54 39.47/39.47 34.23/34.23 40.45/51.55 31.14/30.69 26.27/27.35 21.48/25.76 36.33/35.96
56.48/60.85 39.01/39.01 34.13/34.13 41.97/53.43 31.42/31.86 27.14/28.79 21.42/26.00 35.41/35.17
56.70/61.34 39.12/39.13 33.96/33.96 41.76/53.43 31.51/32.57 27.37/29.18 21.44/26.46 35.87/35.70
Mean 56.48/60.24 39.2/39.2 34.11/34.11 41.39/52.80 31.36/31.71 26.93/28.44 21.45/26.07 35.87/35.61
Variance 0.03/1.12 0.03/0.03 0.01/0.01 0.34/0.59 0.02/0.45 0.17/0.46 0.00/0.06 0.11/0.08
ViT-L-14
Task Unrelated 57.45/59.62 41.81/41.81 37.03/37.03 50.52/56.10 37.13/35.07 26.37/24.35 25.09/29.11 34.96/34.65
56.83/56.52 42.50/42.50 37.51/37.51 51.22/54.32 37.57/31.89 24.56/20.30 23.19/24.47 38.91/38.06
52.46/51.96 41.84/41.84 37.21/37.21 49.65/56.01 37.40/31.53 23.63/19.79 24.99/28.22 39.20/39.63
Mean 55.58/56.03 42.05/42.05 37.25/37.25 50.46/55.48 37.37/32.83 24.85/21.48 24.42/27.27 37.69/37.45
Variance 3.7/7.42 0.08/0.08 0.03/0.03 0.31/0.50 0.02/1.9 0.97/3.12 0.57/3.03 2.81/3.24
Task Related 57.33/61.47 42.58/42.58 37.28/37.28 51.20/55.85 38.00/38.00 27.44/26.16 25.21/29.34 37.90/37.80
57.75/61.86 42.64/42.64 37.41/37.41 51.32/55.85 37.89/38.10 27.44/26.09 24.99/29.08 38.22/38.06
58.11/62.32 42.61/42.61 37.36/37.36 51.29/55.66 38.32/38.57 27.90/26.48 25.24/29.58 38.59/38.32
Mean 57.73/61.88 42.61/42.61 37.35/37.35 51.27/55.79 38.07/38.22 27.59/26.24 25.15/29.33 38.24/38.06
Variance 0.08/0.09 0.00/0.00 0.00/0.00 0.00/0.01 0.02/0.05 0.04/0.02 0.01/0.03 0.06/0.03
@@
  • “a photo of a face with an expression of {class name}.”,
@@
  • “a photo of a face exuding {class name}.”,
  • “a photo of a person embodying {class name}.”,
  • “a good photo capturing someone’s {class name}.”,
  • “a cropped photo emphasizing a face expressing {class name}.”,
@@
As shown in Tab. III, except for one UAR on the DFEW dataset, our method outperforms the CLIP model on all the FER test sets, demonstrating the effectiveness of the proposed method. Notably, our method achieves the best and second-best results consistently under the ViT-L-14 model. In addition, in our method, we extend the length of the feature embedding post our projection head to accommodate the LLM encoder output. To explore the potential impact of feature embedding length on zero-shot results, we also conduct up-sampling on the CLIP feature representations. The results depicted in Table III indicate that longer or sparser feature embeddings from simple up-sampling do not contribute to improved zero-shot results. It should be noted that, compared with the CLIP model, only one extra projection head is added on top of the two encoders in our method and this projection head was trained in an unsupervised manner.
Evaluation of Different Projection Heads: In our method, to align the original CLIP feature representations into task-specific LLMs knowledge space, we propose to utilize a projection head as a linear mapping to fulfil this. To investigate if the improvement of the complexity of the projectors can further improve the alignment, we also conduct experiments with 2-layer MLP, 3-layer MLP, and Transformer Encoders. As shown in Tab. IV, even with a simply learnable matrix and fewer parameters, our strategy is more powerful and efficient. Compared to the Transformer Encoders, except for the WAR results on FERV39k, ours shows much better results than more complex Transformer Encoders. We believe that more complex non-linear models will entail overfitting to the LLMs feature space, performing worse in downstream zero-shot FER tasks. The linear projection head adopted in our method is more suitable* and interpretable* for mapping the original CLIP feature space to a task-specific one, maintaining the capacity of the existing model to capture robust, versatile representations while tailoring it to the FER domain with minimal trainable parameters.
@@
Evaluation of Different Instructions: To verify the effectiveness of the instruction adopted in our method, we conducted experiments on two types of instructions: task-related and task-unrelated. Regarding the task-unrelated instructions, we consider both empty and random text as input and take the average of them as the final results. For the task-related instruction, we provide results with three instructions and their average, including “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”, “Please play the role of a facial expression recognition expert. Describe the facial expression of the person in the image.”, and “Please describe the detailed facial actions of the person in the image.”. As shown in Tab. V, the task-related instructions are superior to task-unrelated instructions on both SFER and DFER, in which there are 7.41%, 12.60%, and 5.85% WAR (accuracy) improvements on RAF-DB (SFER) and 4.06%, 7.40%, and 5.39% WAR (accuracy) improvements on DFEW (DFER). In addition, the three task-related instructions show a small variance in downstream zero-shot prediction. We believe that task-related instructions facilitate LLMs to generate semantic features with task-aware ability, providing better objects for projection head optimisation.
@@
IV-E Visualizations
Moreover, to offer a more extensive comparison with the CLIP model, we include the learned confusion matrices of both our method and the CLIP model, displayed in Fig. 5 and Fig. 6. From the confusion matrix on both static FER and dynamic FER datasets, we can see that our method learned better representations for each expression category, resulting in a more balanced accuracy across all categories. Specifically, our method can better distinguish between neutral expressions and other expressions. This suggests that our method captures more nuanced facial expression features from LLMs.
@@ Acknowledgments
This research utilised Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT. We acknowledge the assistance of the ITS Research team at Queen Mary University of London. Zengqun Zhao is funded by Queen Mary Principal’s PhD Studentships. We thank Ziquan Liu, James Oldfield, and Yimeng Gu for their valuable comments.

### v2

@@ IV-D Ablation Analysis
Evaluation of the Projection Head Effectiveness: Our method is built on the CLIP model [20], to evaluate the effectiveness of the learned projection head, we compare our method to the CLIP model from two views: various ViT models and various zero-shot prompts. Our method uses the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”. As shown in Tab. II, under different ViT models, our method outperforms the plain CLIP model on all the test sets. With ViT-B-32, our method improves the average UAR by 2.71% and average WAR by 4.42% over eight test sets. With ViT-B-16, our method improves the average UAR by 5.47% and average WAR by 6.75% over eight test sets. With ViT-L-14, our method improves the average UAR by 8.07% and average WAR by 8.23% over eight test sets.
Furthermore, we compare our method with CLIP using different prompts, in which both single prompt and prompt ensembles are considered. Specifically, for a single prompt, we utilize “{class name}.”, “an expression of {class name}.”, and “a photo of a face with an expression of {class name}.”, respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations. All the prompt templates are outlined below:
TABLE IV: Zero-shot FER results of our method with different instructions. Both task-related and task-unrelated instructions are investigated. All models use the prompt “a photo of a face with an expression of [class].” for zero-shot prediction.
@@
ViT-B/32
Task Unrelated 33.53/43.42 31.24/31.24 27.18/27.18 30.43/36.69 22.41/23.60 20.28/19.59 16.10/20.41 28.28/29.66
30.49/24.19 27.77/27.78 24.40/24.41 30.54/33.44 20.03/16.19 18.38/15.42 13.21/14.83 29.33/29.66
37.09/36.70 31.23/31.24 26.73/26.73 38.12/36.15 22.26/22.43 18.39/18.07 14.82/17.29 30.65/33.33
Mean 33.70/34.77 30.08/30.09 26.10/26.11 33.03/35.43 21.57/20.74 19.02/17.69 14.71/17.51 29.42/30.88
Variance 7.28/63.49 2.67/2.66 1.48/1.47 12.96/2.02 1.18/10.58 0.80/2.97 1.40/5.21 0.94/2.99
Task Related 39.56/42.14 31.73/31.73 27.63/27.63 37.83/37.04 24.25/25.87 21.48/21.25 17.53/20.27 29.72/31.23
38.63/45.11 31.06/31.07 27.58/27.58 39.46/37.17 25.18/29.48 21.94/24.09 17.96/21.08 30.87/32.81
38.39/44.69 31.35/31.35 28.03/28.03 41.03/36.82 25.39/30.22 22.11/24.61 17.89/21.21 31.61/33.60
Mean 38.86/43.98 31.38/31.38 27.75/27.75 39.44/37.01 24.94/28.52 21.84/23.32 17.79/20.85 30.73/32.55
Variance 0.26/1.73 0.08/0.07 0.04/0.04 1.70/0.02 0.25/3.61 0.07/2.18 0.04/0.17 0.61/0.97
$\Uparrow$ (mean) 5.16/9.21 1.30/1.30 1.64/1.64 6.41/1.58 3.37/7.78 2.83/5.62 3.08/3.34 1.31/1.66
ViT-B/16
Task Unrelated 41.68/41.53 35.01/35.01 30.55/30.56 39.36/51.16 23.48/27.22 22.07/24.20 16.30/19.67 30.65/31.23
37.23/34.75 34.57/34.58 29.73/29.73 36.25/43.13 24.25/28.85 20.61/23.44 17.13/19.68 29.09/30.18
45.61/46.64 39.23/39.24 33.60/33.61 40.49/50.53 26.35/28.54 22.34/23.95 18.15/19.71 30.87/31.76
Mean 41.51/40.97 36.27/36.28 31.29/31.30 38.70/48.27 24.69/28.20 21.67/23.86 17.19/19.69 30.20/31.06
Variance 11.72/23.72 4.41/4.42 2.77/2.78 3.21/13.29 1.47/0.50 0.58/0.10 0.57/0.00 0.63/0.43
Task Related 48.96/54.50 39.98/39.98 34.40/34.40 40.81/53.02 27.12/31.75 24.78/27.99 19.17/23.35 32.84/33.86
48.61/54.14 39.12/39.13 33.63/33.63 39.85/53.94 25.82/31.21 25.28/28.69 18.88/23.62 30.95/32.02
48.94/55.41 39.27/39.27 34.08/34.08 39.91/54.51 26.27/31.97 25.28/29.03 18.90/23.96 30.71/32.02
Mean 48.84/54.68 39.46/39.46 34.04/34.04 40.19/53.82 26.40/31.64 25.11/28.57 18.98/23.64 31.50/32.63
Variance 0.03/0.29 0.14/0.14 0.10/0.10 0.19/0.38 0.29/0.10 0.06/0.19 0.02/0.06 0.91/0.75
$\Uparrow$ (mean) 7.33/13.71 3.19/3.18 2.74/2.74 1.49/5.55 1.71/3.44 3.44/4.71 1.79/3.96 1.30/1.58
ViT-L/14
Task Unrelated 55.83/62.32 41.41/41.41 35.76/35.76 44.86/54.72 36.39/43.82 26.29/31.92 23.94/26.94 37.89/40.16
58.02/62.39 44.78/44.78 38.36/38.36 48.17/54.42 38.80/44.34 26.24/31.11 23.34/27.02 37.92/40.94
58.12/59.06 43.24/43.24 37.49/37.48 47.35/53.33 38.63/42.03 25.05/27.67 23.35/25.44 38.44/39.90
Mean 57.32/61.26 43.14/43.14 37.20/37.20 46.79/54.16 37.94/43.40 25.86/30.23 23.54/26.47 38.08/40.33
Variance 1.12/2.41 1.90/1.90 1.17/1.17 1.98/0.36 1.21/0.98 0.33/3.39 0.08/0.53 0.06/0.20
Task Related 58.70/65.37 44.27/44.27 38.44/38.43 48.28/55.42 40.16/47.09 27.45/31.07 23.95/26.98 38.72/40.42
58.35/66.07 44.52/44.53 38.18/38.18 46.93/55.75 40.73/47.19 28.03/31.62 24.27/27.47 40.00/41.21
57.19/65.74 44.44/44.44 38.11/38.11 47.88/55.50 40.88/47.28 28.23/31.83 24.26/27.37 40.72/41.99
Mean 58.08/65.73 44.41/44.41 38.24/38.24 47.70/55.56 40.59/47.19 27.90/31.51 24.16/27.27 39.81/41.21
Variance 0.42/0.08 0.01/0.01 0.02/0.02 0.32/0.02 0.10/0.01 0.11/0.10 0.02/0.04 0.69/0.41
$\Uparrow$ (mean) 0.76/4.47 1.27/1.27 1.04/1.04 0.90/1.40 2.65/3.79 2.04/1.27 0.62/0.81 1.73/0.87
  • “{class name}.”,
@@
  • “a photo of a face exuding {class name}.”,
@@
  • “a photo of a person embodying {class name}.”,
  • “a good photo capturing someone’s {class name}.”,
@@
  • “a photo of a face with an expression of {class name}.”,
For each prompt, we compare our method with the CLIP model one by one. As shown in Tab. III, except for seven values slightly lower than the CLIP model, our method outperforms the CLIP model on all the rest FER test sets, demonstrating the effectiveness of the proposed method. It should be noted that, compared with the CLIP model, only one extra projection matrix is added on top of the two encoders in our method and this projection matrix was trained in an unsupervised manner.
Evaluation of Different Instructions: To verify the effectiveness of the instruction adopted in our method, we conducted experiments on two types of instructions: task-related and task-unrelated. Regarding the task-unrelated instructions, we consider both empty and random text as input and take the average of them as the final results. For the task-related instruction, we provide results with three instructions and their average, including “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”, “Please play the role of a facial expression recognition expert. Describe the facial expression of the person in the image.”, and “Please describe the detailed facial actions of the person in the image.”.
As shown in Tab. IV, the task-related instructions are superior to task-unrelated instructions on both SFER and DFER, in which there are 3.14% average UAR improvements and 4.02% average WAR improvements over eight test sets with ViT-B-32, 2.87% average UAR improvements and 4.86% average WAR improvements over eight test sets with ViT-B-16, and 1.38% average UAR improvements and 1.87% average WAR improvements over eight test sets with ViT-L-14. In addition, the three task-related instructions show a small variance in downstream zero-shot prediction. We believe that task-related instructions facilitate LLMs to generate semantic features with task-aware ability, providing better objects for projection head optimisation.
TABLE V: Zero-shot FER results of our method with different projection heads. All models are based on ViT-L-14, and use the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.” and the prompt “a photo of a face with an expression of [class].” for zero-shot prediction. 2-MLP and 3-MLP denote 2-layer and 3-layer multilayer perceptron, respectively. TE denotes Transformer encoder. LM denotes our learnable linear matrix.
Projection Heads Static FER (UAR/WAR) Dynamic FER (UAR/WAR)
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
2-MLP 45.05/59.49 35.46/35.47 29.73/29.73 39.66/46.61 31.61/37.35 25.21/26.91 17.62/18.92 29.92/31.76
3-MLP 42.28/61.05 34.95/34.95 30.18/30.18 36.60/51.04 29.67/36.56 25.30/27.25 17.81/19.47 31.91/33.07
TE 34.22/53.13 28.89/28.89 23.88/23.88 33.39/45.20 29.33/37.36 26.06/32.31 17.94/23.20 27.16/28.08
LM 58.70/65.37 44.27/44.27 38.44/38.43 48.28/55.42 40.16/47.09 27.45/31.07 23.95/26.98 38.72/40.42
TABLE VI: Zero-shot FER results of varied image and text encoders. PH denotes the projection head trained in our method.
Image Encoder Text Encoder Static FER Dynamic FER
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
CLIP CLIP 47.22/41.13 34.46/34.47 29.96/29.96 33.82/46.67 32.02/38.12 21.97/28.99 20.42/25.58 35.50/38.32
CLIP+PH LLM 24.03/38.62 24.03/24.04 19.03/19.03 21.86/37.07 21.31/15.31 18.02/18.72 10.10/5.76 19.33/19.16
LLM CLIP+PH 25.81/29.11 19.36/19.35 17.62/17.60 22.59/25.34 13.93/3.77 15.95/10.69 9.66/8.90 19.11/17.06
LLM LLM 19.46/41.23 17.00/17.00 14.88/14.88 18.01/32.23 16.97/22.73 15.18/19.28 10.25/14.66 17.94/19.42
CLIP+PH CLIP+PH 58.70/65.37 44.27/44.27 38.44/38.43 48.28/55.42 40.16/47.09 27.45/31.07 23.95/26.98 38.72/40.42
Figure 3: Comparison of the video-to-text retrieval performance on the subset of the MAFW.
@@
Evaluation of Different Projection Heads: In our method, to align the original CLIP feature representations into task-specific LLM knowledge space, we propose to utilize a projection head as a linear mapping to fulfil this. To investigate if the improvement of the complexity of the projectors can further improve the alignment, we also conduct experiments with 2-layer MLP, 3-layer MLP, and Transformer Encoders.
As shown in Tab. V, even with a simply learnable matrix and fewer parameters, our strategy is more powerful and efficient. Compared to the Transformer Encoders, except for the WAR results on FERV39k, ours shows much better results than more complex Transformer Encoders. We believe that more complex non-linear models will entail overfitting to the LLMs feature space, performing worse in downstream zero-shot FER tasks. The linear projection head adopted in our method is more suitable and interpretable for mapping the original CLIP feature space to a task-specific one, maintaining the capacity of the existing model to capture robust, versatile representations while tailoring it to the FER domain with minimal trainable parameters.
Evaluation of Vision-to-Text-Description Understanding Ability: To verify the vision-to-text-description ability of our method, we conducted experiments using facial videos and their corresponding facial action descriptions. Specifically, we performed a video-text retrieval task on a subset of the MAFW dataset [49]. This subset comprises 8,034 sample-level text descriptions, each detailing specific facial actions. We utilized Precision@k as the metric for video-to-text retrieval and compared the performance of our method against several other vision-language models. As illustrated in Fig. 3, our Exp-CLIP model outperforms the other methods, which means our model retrieves more relevant text descriptions for a given facial video compared to other models and excels in understanding and retrieving complex and detailed facial action descriptions.
Evaluation of Varied Image and Text Encoders: In our method, we proposed to project the original CLIP feature space into a task-aware space, in which the LLM is adopted as a guide. Essentially, on top of the CLIP model, the proposed method is trying to distillate task-specific knowledge from the LLM encoder. To investigate if the LLM can do better than our method on zero-shot FER tasks, we conducted experiments using the LLM as either an image encoder (with I2T module), text encoder, or both. As shown in Tab. VI, adopting the LLM as either an image encoder or text encoder results in a worse zero-shot preference. Even when using the LLM as both an image and text encoder, the results remain undesirable and significantly lower than the plain CLIP model and our method. We believe the LLM excels in the semantic perception of images but performs poorly in recognition tasks. By contrast, based on the CLIP model, our method is more efficient and accurate in zero-shot FER task.
IV-E Visualizations
@@
Moreover, to offer a more extensive comparison with the CLIP model, we include the learned confusion matrices of both our method and the CLIP model, displayed in Fig. 6 and Fig. 7. From the confusion matrix on both static FER and dynamic FER datasets, we can see that our method learned better representations for each expression category, resulting in a more balanced accuracy across all categories. Specifically, our method can better distinguish between neutral expressions and other expressions. This suggests that our method captures more nuanced facial expression features derived from LLMs.
@@ Acknowledgments
This research utilised Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT. We acknowledge the assistance of the ITS Research team at Queen Mary University of London. Zengqun Zhao is funded by Queen Mary Principal’s PhD Studentships. We thank Dr. Ziquan Liu, James Oldfield, and Yimeng Gu for their valuable comments.